2y + y + 2y = g(t), (O) = 0, y(0) = 0 where g) 5 St<20 10, 0<t<5 and t > 20
Answer #1


git) 2y + y +24 410)= 0, yco) so, g(t) =} 5 stC20 6 ostas ft2,20 2y+yt + 2y t ha a OD both side, 2 glt) Taking Laplace tra 205 (25% staj yes) -5s e - s Ss 205 y (s) e e s (25+5+23 LT 205 e e 2 28 - Taking Inverse both side, ss y lt) = s(25+s+23 co/۱ دی 12 2 s(283+8+3) (st/4)² + 15 16 11 -1 5(25+s+2) st teh + 1 4 (s+44) + 2.5 2 2 4 en arte) = ***822-23) [1-ci il * ( 78Equation 2 becomes 200 5 y lt) e e 다. (234 +2) ( - c (2+2+2) | utt-20) att ~ 20) 01 2 -tt-5) e 이 140) + 5 - Lult-s) lt-s) ult